# MCTR: Multi Camera Tracking Transformer

### v1

@@ 1 Introduction
The predominant approach to multi-camera tracking thus far has revolved around heuristic techniques, which amalgamate various components such as person re-identification, single-camera tracking, homography estimation, and clustering. These techniques have demonstrated commendable performance; however, they are inherently heuristic, and their performance often hinges on the effectiveness of hand-crafted rules and the quality of heuristics used. Recently, the application of transformer-based algorithms to tracking has received significant attention [40; 50; 29; 5; 55] due to their ability to encapsulate the entirety of the tracking process within a unified, end-to-end framework. However, this transition toward end-to-end tracking primarily applies to single-camera scenarios. In the domain of multi-camera tracking, particularly when dealing with highly overlapping camera views, there has been little work in this direction.
In this paper, we propose the Multi Camera Tracking tRansformer (MCTR), a novel approach to multi-camera multi-object tracking that adopts an end-to-end architecture to track multiple objects across multiple camera feeds. MCTR builds on recent end-to-end object detection models such as DETR (DEtections TRansformer) [7], adding two new components to facilitate multi-camera tracking: a tracking module and an association module. The flow of MCTR is conceptually simple: the object detectors are independently applied on every camera feed to generate view-specific detections and detection embeddings; the tracking module maintains a set of track embeddings and updates them using the information in the view-specific detection embeddings; and the association module generates an assignment of detections to tracks based on the respective detection and track embeddings. The entire system is trained end-to-end using a new loss formulation. This novel approach offers a more principled solution to multi-camera tracking, by reducing the reliance on heuristic components and integrating the entire tracking process into a coherent, data-driven framework.
@@
As is the case with all transformer-based models, the track embeddings as well as the detection embeddings do not have a particular order, and, thus, do not have a fixed relation with ground truth labels. This makes it challenging to devise a loss for training the model. DETR solves this issue for object detection by first finding the best assignment of detection embeddings to ground truth annotations using the Hungarian algorithm [23], and then computing classification and detection losses based on this assignment. MOTR and other approaches extend this idea to single-camera tracking by forcing each track query that has been previously associated with a ground truth object to keep detecting the same object in future frames. It is unclear, however, how to extend this procedure to keep track assignments consistent not only across time, but also across camera views. Furthermore, this approach is highly dependent on having a good association between detections and ground truth in the first frame, as it can not be changed in subsequent frames. In this paper we take a different approach to devise a training loss that avoids these issues. To calculate view-specific classification and detection losses we follow DETR’s approach and use the Hungarian algorithm in every frame and camera view independently to locally associate detections with ground truth. To maintain global tracking consistency, we use an approach akin to the attention mechanism in scaled dot-product attention to generate probabilistic assignments of local detections to global track embeddings. For pairs of detections in different cameras or different frames we calculate the probability, under the model, that the detections belong to the same track by integrating over the track assignment. We then use a negative log-likelihood loss to train the model to predict a high probability for pairs of detections that are indeed associated with the same ground truth track, and a low probability for pairs of detections that are associated with different ground truth tracks.
@@
Figure 1: Example of multi-camera frame from the MMPTrack dataset with 6 camera angles
@@ 2.1 Single Camera Tracking
Multi-object tracking within a single-camera setup has been extensively studied. Tracking by detection approaches [44; 38; 52; 6; 25] leverages object detectors [11; 14; 34], to identify objects in individual frames and perform track association by employing Kalman Filter [43] for state prediction and Hungarian Matching [23] facilitates track correspondence between frames. Recently, end to end tracking methods such as [40; 50; 5; 54; 53; 49; 55; 29] have emerged, extending query based object detection for tracking [7]. For example, MOTR [50], MOTRv2 [53] and TrackFormer [29] propagate track queries across frames, iteratively updating them with image features for long-term detection and tracking. TransTrack [40] and P3Aformer [54] utilize a location-based cost matrix for bipartite matching [23]. In contrast to tracking by detection methods, end-to-end tracking methods are data-driven and avoid dependence on handcrafted heuristics. Despite their effectiveness, single-camera tracking encounters challenges, especially in cluttered and crowded environments with occlusions. The reliance on a single viewpoint limits the system’s robustness in complex scenarios. To alleviate this, we extend the tracking by query propagation paradigm to the more intricate task of multi-camera multi-object tracking.
@@ 2.2 Multi-Camera Tracking
The multi-camera tracking (MCT) domain has witnessed diverse methodologies to tackle challenges in complex surveillance environments. A prevalent strategy involves a distributed approach, where single-camera tracking precedes hierarchical clustering [30], non-negative matrix factorization (NMF) [42], or other merging and association algorithms for tracklets [4; 24; 33; 17; 20; 39; 46; 8; 22; 45; 36]. Alternatively, global or centralized methods bypass single-camera tracking, focusing on detecting individuals in each camera view and subsequently globally associating detections into tracklets [47; 27]. Occupancy maps derived from multi-camera views have been explored [12; 48], and trajectory prediction is employed in certain works, with variations in global or distributed approaches for multi-view association [47; 19]. Iguernaissi et al. [18] provides a comprehensive survey of the MCT papers.
Notable contributions in MCT include TRACTA [17], utilizing RNMF for cross-camera tracklet matching, and DMCT [48], incorporating a perspective-aware GroundPoint Network, occupancy heatmap estimation, and a glimpse network for person detection and tracking. Recently, ReST [9] proposed a two-stage association method using a reconfigurable graph model. In the vision-based autonomous driving systems domain, efforts focus on multi-view 3D object tracking, building upon end-to-end multi-view 3D object detection methodologies [10; 41]. Works such as MUTR3D [51], PF-Track [32], ViP3D [15], and DQTrack [26] leverage a 3D track query-based approach for coherent object tracking across multiple cameras and frames. DQTrack [26] introduces a decoupled-query paradigm for camera-based 3D multi-object tracking, and ViP3D [15] pioneers a fully differentiable trajectory prediction approach. PF-Track [32] emphasizes spatio-temporal continuity.
In our approach, we conduct global track association for detections in each individual view, departing from single-camera tracklet-to-tracklet association. Diverging from existing global association methods, our fully end-to-end trainable framework based on transformers eschews heuristics or handcrafted pipelines for data-driven learning. Aligned with 3D object tracking methodologies, our track association loss is entirely differentiable and trained end-to-end. In contrast to 3D object tracking in autonomous driving, our work focuses on 2D person tracking across overlapping cameras.
@@ 3 Method
The system consists of a detection module, a tracking module, and an association module. The detection module operates independently for each camera view and generates a set detection embeddings. These embeddings encapsulate essential information regarding object detections within each individual view and are used to generate a bounding box prediction, a class prediction, and to inform the downstream modules. The tracking module maintains a set of track embeddings. The track embeddings contain global information about each object and are used to maintain consistent and coherent object identities across camera views and time. The track embeddings are updated at each frame using the information in the detection embeddings from all the camera views. The association module is tasked with producing a probabilistic assignment of detections to identities based on the information in the respective track and detection embeddings.
@@
3.1 Detection Module
In this paper we employ a vanilla DETR (DETection TRansfomer) for the detection module, but any other end-to-end detector architecture can be used. A DETR model operates independently on each camera view. Briefly, the architecture of the DETR model includes a backbone to extract image features and a transformer encoder and decoder. The spatial features at the output of the backbone are added to a positional encoding and input to the encoder transformer which performs self attention and outputs a set of spatial features. The DETR transformer decoder takes a set of detection query embeddings as input and outputs a set of detection embeddings. Each of the output embeddings are then passed through an MLP to predict the object class (or a "no object" class) and its bounding box. An overview of DETR is shown in figure 3. For more details see [7].
@@ 3.2 Tracking Module
The track embeddings are initialised at the first frame to learned "track query embeddings". Like the "query inputs" to DETR, some of the embeddings will take on the role of representing one identity over time and others will remain unassigned. When a track embedding remains unassigned to a detection after a certain time, it is reset to it’s initial embedding.To update the track embeddings with detections from each camera, the cross-attention section has a set of multi-head cross-attention modules tailored to each camera view, with each module computing cross-attention between the current track embeddings and the detection embeddings from the corresponding view. The track embeddings act as "queries", and detection information as keys and values. Because the relationship between objects and tracks depends on the camera position, each cross-attention module has its own distinct parameters (i.e. there is no parameter sharing between the cross-attention modules corresponding to the different views). The outputs of all the view specific cross-attention modules are averaged and passed through self-attention and feed-forward layers to obtain updated track embeddings.
The self-attention module in the tracking module is intended to introduce competition between the track embeddings. If one track query ’claims’ a particular identity strongly, other track queries can "observe" this and are encouraged not to take on the same identity through the attention mechanism. This is similar to the way self attention works in the decoder for DETR to discourage 2 query embeddings from claiming the same output detection.
@@ 3.4 Training Loss
$\mathcal{L}_{across\_cams}=-\frac{1}{N_{p}}\sum_{d_{1},d_{2}}\bigg{(}$ $y_{st}(d_{1},d_{2})\cdot\log(P_{st}(d_{1},d_{2}))$
$+(1-y_{st}(d_{1},d_{2}))\cdot\log(1-P_{st}(d_{1},d_{2}))\bigg{)}$
@@
$\mathcal{L}=\mathcal{L}_{det}+\mathcal{L}_{det\_aux}+\mathcal{L}_{across\_cams}+\mathcal{L}_{across\_frames}+\mathcal{L}_{track\_IOU}+\mathcal{L}_{track\_GIOU}$
@@ 3.5 Training Protocol
Figure 6: Effect of training protocol.
Figure 7: Inference time compute resources.
@@ 3.6 Inference
Figure 7 shows the memory and time MCTR uses for inference as a function of the number of cameras. Both time and memory scale linearly with the number of camera. With 1 camera, the model has about 44M parameters and takes 233Mb of memory and can run at 21 FPS on a single GeForce RTX 2080 Ti GPU. Each additional camera adds 1M parameters, 68Mb of memory and reduces the FPS by about 2. Thus MCTR would be well suited for real-time multi-camera tracking applications.
@@ 4 Experimental Results
We contrast MCTR with the single-camera end-to-end tracker MOTR [50], when the later is applied to each camera view independently.  [^1] While the performance of MOTR may be improved by using pre-trained object detectors to generate detection queries, as in MOTR v2 [53], the same techniques are equally applicable to MCTR. In fact MCTR can be used directly with detectors such as YOLO foregoing DETR altogether. However, this would distract from the goal of understanding the behaviour of end-to-end approaches. We also attempted to compare with ReST [9], a recent multi-stage multi-camera tracking approach. Unfortunately, despite significant efforts, we were unable to get ReST to work on our evaluation datasets. This highlights the shortcomings of multi-stage methods that require significant effort to ensure that all of the heuristic components are well tuned for the dataset. In contrast end-to-end methods require much less manual tuning, learning to adapt to the characteristics of each dataset.
@@
The models are evaluated using popular metrics in single-camera multi-object tracking literature: HOTA[28], IDF1[37] and MOTA[3]. The metrics are computed independently for each camera view and test clip, and the results are averaged to get the final performance for each environment. This is similar to the protocol used in the MMP-Tracking challenge [1]. For MCTR we calculate the Inter-Camera Association Accuracy (ICAA), which measures the fraction of pairs of detections from different views at the same time frame that the model correctly associates with same ground track ID.
@@ 4.1 MMPTrack dataset
One artifact of the way the MMPTrack annotations have been collected is that ground-truth bounding boxes are given for all views, even if the person might be fully occluded in a particular view. In most environments, occlusions are rare, making it less problematic. However, in retail setting, a large fraction of people are occluded by shelves in every view. Training a detector with these annotations would be ambiguous and lead to lower performance. To mitigate this issue, we filter out annotations of occluded persons: we employ a pre-trained YOLOX model [13] to detect people in each frame, and ground truth annotations without an IoU greater than 0.01 with a YOLOX detection are filtered out. While this approach may eliminate annotations for visible but undetected people by YOLOX, it serves as a reasonable approximation. The models are evaluated on the original labels, so low recall is expected in this environment.
Table 1: Results on MMPTrack dataset
HOTA IDF1 MOTA ICAA
Env. MCTR MOTR MCTR MOTR MCTR MOTR MCTR
Industry 71.81 53.94 80.21 52.10 91.19 89.42 99.22
Lobby 65.33 60.74 68.07 57.55 88.39 84.38 96.01
Cafe 45.25 33.59 49.71 28.94 84.22 74.27 92.78
Office 52.20 44.73 55.01 41.62 88.05 82.79 96.75
Retail 44.25 25.99 58.48 27.80 53.21 48.94 98.87
Retail-TB 62.06 - 71.04 - 75.41 - 98.87
The results of executing MCTR and MOTR on the validation clips from the five MMPTrack environments are presented in Table 1. These results represent averages across various cameras and clips. Across all environments, MCTR demonstrates superior performance compared to MOTR across all metrics. This underscores the effectiveness of principled end-to-end methodologies for multi-camera tracking. The higher IDF1 and HOTA scores indicate that MCTR is able leverage cues from multiple views to address occlusions and maintain more consistent long term tracks. The inter-camera association accuracy for MCTR is consistently above 92% indicating that the model accurately tracks identities across camera views.
@@
As expected, the performance on the retail environment is lower because of the large number of ground-truth detections that are occluded in any particular camera view. While a person might be occluded in one or more camera views, it is usually visible in at least one camera allowing MCTR the opportunity to detect that person. As discussed in section 3.4, MCTR is trained to predict bounding boxes for all views based on the global track embeddings. While these bounding boxes are primarily utilized for computing an auxiliary loss to enhance model training, they can be also used to predict bounding boxes for occluded people that can not be detected by the view-specific detector. When using the track-level bounding boxes for prediction (denoted as Retail-TB in table 1), the performance improves significantly, mainly driven by an increase in recall from 54.28 to 88.73. This is remarkable, because the model has not seen any ground truth annotations in the occluded regions during training. This suggests that the model has learned an approximate geometry and is able to extrapolate it to regions where it has not seen any training data.
@@
To the best of our knowledge, state of the art performance on the MMPTrack dataset has been achieved during the MMP-Tracking challenge[1]. The highest MOTA score on the leaderboard is 87, averaged across all five environments, and the highest IDF1 score is 88.18 [2]. Besides using a plethora of engineered heuristics, both these methods are offline (i.e. they can use information from the future), and make heavy use of camera calibration parameters. MCTR is online, and does not use camera parameters. With a combined MOTA of 81.04 and a combined IDF1 of 62.16, MCTR would rank thrid in MOTA and fifth in IDF1.222Our performance is measured on the validation set as we do not have access to the test set. We assume that validation and test performances are comparable. While the MOTA performance is respectable, there is a large gap in the IDF1 performance. This indicates that maintaining consistent tracking over a long period of time is still challenging for an end-to-end method.

### v2

@@ 1 Introduction
The predominant approach to multi-camera tracking thus far has revolved around heuristic techniques, which amalgamate various components such as person re-identification, single-camera tracking, homography estimation, and clustering. These techniques have demonstrated commendable performance; however, they are inherently heuristic, and their performance often hinges on the effectiveness of hand-crafted rules and the quality of heuristics used. Recently, the application of transformer-based algorithms to tracking has received significant attention [42, 52, 31, 5, 57] due to their ability to encapsulate the entirety of the tracking process within a unified, end-to-end framework. However, this transition toward end-to-end tracking primarily applies to single-camera scenarios. In the domain of multi-camera tracking, particularly when dealing with highly overlapping camera views, there has been little work in this direction.
In this paper, we propose the Multi Camera Tracking tRansformer (MCTR), a novel approach to multi-camera multi-object tracking that adopts an end-to-end architecture to track multiple objects across multiple camera feeds. MCTR builds on end-to-end object detection models such as DETR (DEtections TRansformer) [7], adding two new components to facilitate multi-camera tracking: a tracking module and an association module. The flow of MCTR is conceptually simple: the object detectors are independently applied on every camera feed to generate view-specific detections and detection embeddings; the tracking module maintains a set of track embeddings and updates them using the information in the view-specific detection embeddings; and the association module generates an assignment of detections to tracks based on the respective detection and track embeddings. The model is trained end-to-end using a new loss formulation. This approach offers a more principled solution to multi-camera tracking, reducing the reliance on heuristic components and integrating the entire tracking process into a coherent, data-driven framework.
@@
As is the case with all transformer-based models, the track embeddings as well as the detection embeddings do not have a particular order, and, thus, do not have a fixed relation with ground truth labels. This makes it challenging to devise a loss for training the model. DETR solves this issue for object detection by first finding the best assignment of detection embeddings to ground truth annotations using the Hungarian algorithm [25], and then computing classification and detection losses based on this assignment. MOTR and other approaches extend this idea to single-camera tracking by forcing each track query that has been previously associated with a ground truth object to keep detecting the same object in future frames. It is unclear, however, how to extend this procedure to keep track assignments consistent not only across time, but also across camera views. Furthermore, this approach is highly dependent on having a good association between detections and ground truth in the first frame, as it can not be changed in subsequent frames. In this paper we take a different approach to devise a training loss that avoids these issues. To calculate view-specific classification and detection losses we follow DETR’s approach and use the Hungarian algorithm in every frame and camera view independently to locally associate detections with ground truth. To maintain global tracking consistency, we use an approach akin to the attention mechanism in scaled dot-product attention to generate probabilistic assignments of local detections to global track embeddings. For pairs of detections in different cameras or different frames we calculate the probability, under the model, that the detections belong to the same track by integrating over the track assignment. We then use a negative log-likelihood loss to train the model to assign high probabilities to detection pairs from the same ground truth track and low probabilities to pairs from different tracks.
@@
Figure 1: Example of multi-camera frame from the MMPTrack dataset with 6 camera angles.
@@ 2.1 Single Camera Tracking
Multi-object tracking within a single-camera setup has been extensively studied. Tracking by detection approaches [46, 40, 54, 6, 27] utilize object detectors [11, 15, 36], to identify objects in individual frames and perform track association using Kalman Filter [45] or Hungarian Matching [25]. Recently, end to end tracking methods such as [42, 52, 5, 56, 55, 51, 57, 31] have emerged, extending query based object detection for tracking [7]. For example, MOTR [52], MOTRv2 [55] and TrackFormer [31] propagate track queries across frames, iteratively updating them with image features for long-term detection and tracking. TransTrack [42] and P3Aformer [56] utilize a location-based cost matrix for bipartite matching [25]. In contrast to tracking by detection methods, end-to-end tracking methods are data-driven and avoid dependence on handcrafted heuristics. Despite their effectiveness, single-camera tracking encounters challenges, especially in cluttered and crowded environments with occlusions. The reliance on a single viewpoint limits the system’s robustness in complex scenarios. To alleviate this, we extend the tracking by query propagation paradigm for multi-camera multi-object tracking.
@@ 2.2 Multi-Camera Tracking
The multi-camera tracking (MCT) domain has witnessed diverse methodologies to tackle challenges in complex surveillance environments. A prevalent strategy involves a distributed approach, where single-camera tracking precedes hierarchical clustering [32], non-negative matrix factorization (NMF) [44], or other merging and association algorithms for tracklets [4, 26, 35, 20, 23, 41, 48, 8, 24, 47, 39]. Alternatively, global or centralized methods bypass single-camera tracking, focusing on detecting individuals in each camera view and subsequently globally associating detections into tracklets [49, 29]. Occupancy maps derived from multi-camera views have been explored [12, 50], and trajectory prediction is employed in certain works, with variations in global or distributed approaches for multi-view association [49, 22]. Iguernaissi et al. [21] provides a comprehensive survey of the MCT papers.
Notable contributions in MCT include TRACTA [20], utilizing RNMF for cross-camera tracklet matching, and DMCT [50], incorporating a perspective-aware GroundPoint Network, occupancy heatmap estimation, and a glimpse network for person detection and tracking. Recently, ReST [9] proposed a two-stage association method using a reconfigurable graph model. In the vision-based autonomous driving systems domain, efforts focus on multi-view 3D object tracking, building upon end-to-end multi-view 3D object detection methodologies [10, 43]. Works such as MUTR3D [53], PF-Track [34], ViP3D [16], and DQTrack [28] utilizes a 3D track query-based approach for coherent object tracking across multiple cameras and frames. DQTrack [28] introduces a decoupled-query paradigm for camera-based 3D multi-object tracking, and ViP3D [16] pioneers a fully differentiable trajectory prediction approach. PF-Track [34] emphasizes spatio-temporal continuity.
In our approach, we conduct global track association for detections in each individual view, departing from single-camera tracklet-to-tracklet association. Diverging from existing global association methods, our fully end-to-end trainable framework based on transformers avoids heuristics or handcrafted pipelines for data-driven learning. Our track association loss is fully differentiable and trained end-to-end, similar to 3D object tracking. However, we focus on 2D person tracking across overlapping cameras.
@@ 3 Method
The system consists of a detection module, a tracking module, and an association module. The detection module operates independently for each camera view and generates a set of detection embeddings. These embeddings encapsulate essential information regarding object detections within each individual view and are used to generate a bounding box prediction, a class prediction, and to inform the downstream modules. The tracking module maintains a set of track embeddings. The track embeddings contain global information about each object and are used to maintain consistent and coherent object identities across camera views and time. The track embeddings are updated at each frame using the information in the detection embeddings from all the camera views. The association module is tasked with producing a probabilistic assignment of detections to identities based on the information in the respective track and detection embeddings.
@@ 3.1 Detection Module
In this paper we employ a vanilla DETR (DETection TRansfomer) for the detection module, but any other end-to-end detector architecture can be used. A DETR model operates independently on each camera view. Briefly, the architecture of the DETR model includes a backbone to extract image features and a transformer encoder and decoder. The spatial features at the output of the backbone are added to a positional encoding and input to the encoder transformer which performs self attention and outputs a set of spatial features. The DETR transformer decoder takes a set of detection query embeddings as input and outputs a set of detection embeddings. Each of the output embeddings are then passed through an MLP to predict the object class (or a ”no object” class) and its bounding box. An overview of DETR is shown in figure 3. For more details see [7].
@@ 3.2 Tracking Module
The track embeddings are initialised at the first frame to learned ”track query embeddings”. Like the ”query inputs” to DETR, some of the embeddings will take on the role of representing one identity over time and others will remain unassigned. When a track embedding remains unassigned to a detection after a certain time, it is reset to it’s initial embedding.To update the track embeddings with detections from each camera, the cross-attention section has a set of multi-head cross-attention modules tailored to each camera view, with each module computing cross-attention between the current track embeddings and the detection embeddings from the corresponding view. The track embeddings act as ”queries”, and detection information as ”keys” and ”values”. Because the relationship between objects and tracks depends on the camera position, each cross-attention module has its own distinct parameters (i.e. there is no parameter sharing between the cross-attention modules corresponding to the different views). The outputs of all the view specific cross-attention modules are averaged and passed through self-attention and feed-forward layers to obtain updated track embeddings.
The self-attention module in the tracking module is intended to introduce competition between the track embeddings. If one track query ’claims’ a particular identity strongly, other track queries can ”observe” this and are encouraged not to take on the same identity through the attention mechanism. This is similar to the way self attention works in the decoder for DETR to discourage 2 query embeddings from claiming the same output detection.
@@ 3.4 Training Loss
$\mathcal{L}_{across\_cams}=$ $-\frac{1}{N_{p}}\sum_{d_{1},d_{2}}\left(y_{st}(d_{1},d_{2})\cdot\log(P_{st}(d_{1},d_{2}))\right.$
$+\left.(1-y_{st}(d_{1},d_{2}))\cdot\log(1-P_{st}(d_{1},d_{2}))\right)$
@@
$\mathcal{L}=\,$ $\mathcal{L}_{det}+\mathcal{L}_{det\_aux}+\mathcal{L}_{across\_cams}+\mathcal{L}_{across\_frames}$
$+\mathcal{L}_{track\_IOU}+\mathcal{L}_{track\_GIOU}$
@@ 3.6 Inference
Figure 7: Inference time compute resources.
Figure 7 shows the memory and time MCTR uses for inference as a function of the number of cameras. Both inference time and memory scale linearly with the number of cameras. With 1 camera, the model has about 44M parameters and takes 233Mb of memory and can run at 21 FPS on a single GeForce RTX 2080 Ti GPU. Each additional camera adds 1M parameters, 68Mb of memory and reduces the FPS by about 2. Thus MCTR would be well suited for real-time multi-camera tracking applications.
@@ 4 Experimental Results
We contrast MCTR with the single-camera end-to-end tracker MOTR [52], when the later is applied to each camera view independently.  [^1] While the performance of MOTR may be improved by using pre-trained object detectors to generate detection queries, as in MOTR v2 [55], the same techniques are equally applicable to MCTR. In fact MCTR can be used directly with detectors such as YOLO foregoing DETR altogether. However, this would distract from the goal of understanding the behaviour of end-to-end approaches. We also compare with ReST [9], a recent multi-stage multi-camera tracking approach. When using a pre-trained YOLO detector ReST is unable to maintain track identities over longer time frames leading to very low performance on our evaluation datasets. This highlights the shortcomings of multi-stage methods that require significant effort to ensure that all of the heuristic components are well tuned for the dataset. In contrast, end-to-end methods require much less manual tuning, learning to adapt to the characteristics of each dataset. To gauge tracking performance of ReST, with an ideal detector, we also report results for ReST when using the ground truth bounding boxes as detections (ReST-GT).
Figure 8: Results on MMPTrack dataset.
@@
The models are evaluated using popular metrics in single-camera multi-object tracking literature: HOTA[30], IDF1[38] and MOTA[3]. The metrics are computed independently for each camera view and test clip, and the results are averaged to get the final performance for each environment. This is similar to the protocol used in the MMP-Tracking challenge [2]. For MCTR and ReST we evaluate the cross-view association performance using the AIDF1 and MOAA metrics as proposed in [18, 17, 13]. AIDF1 is the association $F_{1}$ score calculated as the geometric mean of precision and recall of predicted pairwise object associations averaged across all frames and all pairs of camera views. MOAA (Multi Object Association Accuracy) follows the calculation of MOTA, but ID switches are measured across all pairs of views rather than across frames.
@@ 4.1 MMPTrack dataset
One artifact of the way the MMPTrack annotations have been collected is that ground-truth bounding boxes are given for all views, even if the person might be fully occluded in a particular view. In most environments occlusions are rare, making it less problematic. In the retail setting, however, a large fraction of people are occluded by shelves. Training a detector with these annotations would be ambiguous and lead to lower performance. To mitigate this issue, we filter out annotations of occluded persons: we employ a pre-trained YOLOX model [14] to detect people in each frame, and ground truth annotations with a lower IOU than 0.01 with a YOLOX detection are filtered out. While this approach may eliminate annotations for visible people that are undetected by YOLOX, it serves as a reasonable approximation. The models are evaluated on the original labels, so low recall is expected in this environment.
Figure 8 shows the performance of MCTR, MOTR, ReST and ReST-GT ( ReST with ground truth bounding boxes) on the validation clips from the five MMPTrack environments  [^2] ReST-GT inference did not finish on the Retail environment, and ReST did not finish on Lobby environment. These results represent averages across various cameras and clips. MCTR outperforms ReST and MOTR across all environments and metrics, showcasing the effectiveness of end-to-end methods for multi-camera tracking. Its higher IDF1 and HOTA scores highlight MCTR’s ability to use multi-view cues to handle occlusions and maintain consistent long-term tracks. Compared to ReST-GT, MCTR performs better on the HOTA and IDF1 metrics, showing that even with a perfect detector, ReST temporal tracking performance is lower than MOTR. On MOTA, MOAA and AIDF1, which weight detection accuracy more heavily, ReST-GT performs better than MCTR due to its perfect detector.
Figure 9: Results on AI City Challenge dataset.
@@
As mentioned above, in the retail environment there are a large number of ground-truth detections that are occluded in any particular camera view resulting in a significant number of missed detections. While a person might be occluded in one or more camera views, it is usually visible in at least one camera allowing MCTR the opportunity to detect that person. As discussed in section 3.4 MCTR is trained to predict bounding boxes for all views based on the global track embeddings. While these bounding boxes are primarily utilized for computing an auxiliary loss to enhance model training, they can be also used to predict bounding boxes for occluded people that can not be detected by the view-specific detector. When using the track-level bounding boxes for prediction (denoted as MCTR-TB in figure 8), the performance improves significantly on the retail environment, mainly driven by an increase in recall from 54.28 to 88.73. This is remarkable, because the model has not seen any ground truth annotations in the occluded regions during training. This suggests that the model has learned an approximate geometry and is able to extrapolate it to regions where it has not seen any training data. On the other environments, the HOTA and IDF1 performance remains similar, but there is a decrease in the MOTA score due to a somewhat worse detection bounding box accuracy.
Besides the ability to predict the location of occluded people, the auxiliary losses described in section 3.4 are instrumental in improving the long term tracking performance of MCTR. Table 1 shows the performance, on the industry environment, of MCTR with and without using the auxiliary loss. Training the model to predict detection bounding boxes from the global track embeddings leads to significant improvements in the tracking focussed HOTA and IDF1 scores, with a slight decrease in the more detection focused metrics.
Table 1: Performance of MCTR with and without using auxiliary losses for training
Model HOTA IDF1 MOTA AIDF1 MOAA
MCTR 71.81 80.21 91.19 94.52 90.39
MCTR-noaux 68.00 75.26 92.11 95.00 91.30
Table 2: Comparison with top methods in the MMP-Tracking challenge, camera view evaluation. The ranks MCTR would have achieved for each metric are shown in parentheses.
Model IDF1 MOTA Online Camera parameters FPS
Hikvision 86.36 87 No Yes -
Alibaba 88.18 78 No Yes $<1$
MCTR 62.30 (5) 81.01 (3) Yes No 9
@@
To the best of our knowledge, state of the art performance on the MMPTrack dataset has been achieved during the MMP-Tracking challenge[2]. The highest MOTA score on the leaderboard is 87, averaged across all five environments, and the highest IDF1 score is 88.18 [1]. Besides using a plethora of engineered heuristics, both these methods are offline (i.e. they can use information from the future), and make heavy use of camera calibration parameters. MCTR is online, and does not use camera parameters. With a combined MOTA of 81.01 and a combined IDF1 of 62.30, MCTR would rank thrid in MOTA and fifth in IDF1.333Our performance is measured on the validation set as we do not have access to the test set. We assume that validation and test performances are comparable. While the MOTA performance is respectable, there is a large gap in the IDF1 performance. This indicates that maintaining consistent tracking over a long period of time is still challenging for an end-to-end method.